OpenStudy (anonymous):

sin(pi/2)co(pi/7)+cos(pi/2)sin(pi/7) Can anybody help me please?

OpenStudy (anonymous):

I'm guessing this is about goniometrics, so pi = ?

OpenStudy (anonymous):

It's write the expression as either sine, cosine, and or tangent of a single angle...and the very last bloody one on my worksheet thank the gods -.-

OpenStudy (anonymous):

Hehe, well, first off, pi = 180°, right? Fill that out first

OpenStudy (anonymous):

sin(pi/2) = sin(180/2) = sin 90° ...

OpenStudy (anonymous):

Right, do I need to get that wheel that my teacher keeps telling me to google out? Ummm I totally forgot the name...it's too early for this

OpenStudy (anonymous):

sinAcosB+cosAsinB = sin(A+B) =sin(pi/2 + pi/7)

OpenStudy (anonymous):

sum and difference identity

OpenStudy (anonymous):

simplify parenthesis maybe

OpenStudy (anonymous):

=sin(9pi/14)

OpenStudy (anonymous):

I've never seen that exact wheel before Miel...I'm trying to find the one I book marked. Ah and ehuman...how did you get the 9? I know you multiplied to get 14 but the 9? Did you add?

OpenStudy (anonymous):